OpenStudy (sammyalabamy):

Find the exact solution to the equation. 5^(6-2x)=25 I know I need to use log, but can I have help with setting up the solution?

hartnn (hartnn):

You can either use logs or just compare the exponents, 1. Using log: \(\Large 5^{6-2x} = 25 \\ \log 5^{6-2x} = \log 5^2 \\ (6-2x)\log 5 =2 \log 5 \\ 6-2x = 2\\ ... \) 2. just compare the exponents, \(\Large 5^{6-2x} = 25 \\\Large 5^{6-2x} = 5^2 \\ \implies 6-2x =2 \\ ...\)

hartnn (hartnn):

that is if you still needed help :)

OpenStudy (sammyalabamy):

thank you.